# Algebroid functions on complex manifolds

### v5

@@ 2.2. Multi-leaf complex manifolds of algebroid functions
In this subsection, we will show how to construct a class of complex manifolds using algebroid functions. Such a class of manifolds have a remarkable feature of multifoliation. What’s important is that these multi-leaf complex manifolds are a way to the uniformization of algebroid functions. First of all, we deal with a simple situation, i.e., we equip with a complex structure such that is a non-compact complex manifold.
As a matter of convenience, a multi-valued function element is also called the multi-leaf function element; a (non-holomorphic) meromorphic function element is also called the polar function element; the polar function elements and multi-leaf function elements are called the singular function elements.
Let be an arbitrary function element. Let us consider a neighborhood of which is defined by such a set of all such that is a direct analytic continuation of satisfying i.e., we have Choose a proper if necessary, we may have a biholomorphic mapping such that in which is the unit ball centered at 0 in Define a natural mapping
@@
for some two single-valued components of respectively, where is the branching set of for Now, put the multi-leaf function elements in Set
@@
If then it corresponds to meromorphic function elements in In particular, if then reduce to regular function elements. If then it can correspond to multi-leaf function elements in with where can be separated into single-valued components.
@@
Let be single-valued connected components of Define a mapping
by
where It is easy to check that is a homeomorphism and its inverse is given by
@@
We will show that is a complex atlas of Let It is sufficient to prove the transition mapping
is a biholomorphic mapping. This is automatically a homeomorphism. Via a similar discussion, one only needs to deal with the situation where is a -leaf singular function element and is just a regular function element. For each there exists a single-valued component of such that in some neighborhood of i.e., It is clear that since all the function elements in are meromorphic. A direct computation leads to
@@
Next, we consider the singular part Let be the natural projection by mapping a function element into Then, one sees that is a generic -sheeted analytic covering and
Since is an analytic set of complex dimension not greater than then the complex structure of can be viewed as a natural extension of the complex structure of This implies that we can naturally extend the complex structure of to via Therefore, we obtain:
@@
is a non-compact -leaf complex manifold of complex dimension called a complex manifold of a -valued algebroid function on
@@ 2.3. Zero and pole divisors of algebroid functions
Let be a -valued algebroid function on defined by equation As we have known in Section 2.1, is a pole of if and only if is a zero of However, it provides nothing about the relationship between the multiplicities of a pole of and a zero of The following theorem shows that they agree. Let be single-valued meromorphic components of As a matter of convenience, we write
@@
Let be a zero of Since is an empty set or an analytic set of complex codimension 2, then we have in the sense of divisors, and so we may assume that It yields that
@@
Combined this with (4), we see that reaches the maximal order of a pole of Since then we obtain
It follows that ∎
@@
3. Uniformization of algebroid functions on Hermitian manifolds
This section works on the uniformization of algebroid functions on a non-compact Hermitian manifold. As shown in Section 2.2, we have constructed a multi-leaf complex manifold of a multi-valued algebroid function on a non-compact Hermitian manifold In fact, this realizes the uniformization of by liftting to a meromorphic function on via the natural projection such that In what follows, we shall provide another way towards the uniformization of algebroid functions without constructing multi-leaf complex manifolds. Our idea is motivated by the work of W. Stoll [50] on algebroid reduction.
@@ 3.2. Uniformization problems
which yields that
If then (6) defines an algebraic equation
@@
By this with (5)
To conclude, we obtain
@@
Set with a basis Since (8), then we can write
@@
in the sense of isomorphisms. On the other hand, Lemma 3.6 says that there exist a non-compact complex manifold of complex dimension a proper surjective holomorphic mapping of (generic) sheet number and a holomorphic mapping which separates the fibers of such that is the algebroid reduction of via Hence, we have
@@
Treat the -leaf complex manifold of a -valued algebroid function on we have the natural projection of (generic) sheet number via which is a proper surjective holomorphic mapping.
@@
Let be a generic -sheeted analytic covering of a non-compact Hermitian manifold with branching set Denote by (resp. ) the field of all meromorphic functions on (resp. ).
We give the algebraic relation between and as follows.
Corollary 3.9.
is a -degree algebraic extension of
Proof.
For a meromorphic function on we see that is at most a -valued algebroid function with branching set on So, there exist meromorphic functions on such that
It implies that is a finite algebraic extension of of degree ∎
4. Nevanlinna theory of meromorphic functions on Kähler manifolds
Let be a non-compact complete Kähler manifold of complex dimension with Laplace-Beltrami operator associated to Kähler metric Recall that the Kähler form of is defined by
in a local holomorphic coordinate Moreover, we define the Ricci form of by
Invoking Wirtinger’s formula, we see that the Riemannian volume element of is
@@ 4.1. First Main Theorem
where is the Dirac function on The harmonic measure on with respect to is defined by
@@
Next, we introduce Nevanlinna’s functions. Let be a meormorphic function. Denote by the Fubini-Study form on where is the homogeneous coordinate of and here
so that
@@
We need the following Dynkin formula:
Lemma 4.1 (Dynkin formula, [3, 16]).
@@ 4.2. Second Main Theorem for non-negative Ricci curvature
In the subsection, we assume that has non-negative Ricci curvature and maximal volume growth, i.e.,
where is the volume of geodesic ball centered at a fixed reference point with radius Denote by the Riemannian distance function of from and by the area of which is the boundary of A global Green function of for with a pole at means that
where is the Dirac function on Put
By Li-Tam-Wang [26], there is a number independent of such that
as
If then there exists (see, e.g., [60, 61]) the unique minimal positive global Green function of for with a pole at According to the estimate due to Li-Yau [27] and (10), there exists a constant depending only on such that
Colding-Minicozzi II [11] (see Li-Tam-Wang [26] also) proved the following asymptotic behavior of :
@@
Define the relatively compact -domain for by
Then, exhausts whenever is an increasing sequence with as Moreover, the boundary of is a submanifold of for almost every due to Sard’s theorem. Set
It is evident that gives the positive Green function of for with a pole at satisfying Dirichlet boundary condition. Also, denote by the harmonic measure on with respect to
If then curvature condition implies that is a parabolic Riemann surface, so each global Green function for must change sign. Again, since has maximal volume growth, we see that is conformally equivalent to i.e., there exists a biholomorphic mapping such that We equip with the pull-back metric induced from the standard Euclidean metric on Since gives a global Green function of for with a pole at then it corresponds to a global Green function of for with a pole at :
In this case, we define the relatively compact -domain for by
It is clear that
gives the Green function of for with a pole at satisfying Dirichlet boundary condition. A direct computation gives that the harmonic measure on with respect to is
where is the Riemannian volume element of
Let be a meromorphic function. Take in Section 4.1, then we have the Nevanlinna’s functions of :
4.2.1. Estimates of harmonic measures
We shall use the asymptotic behavior of (see Lemma 4.3) to derive both lower and upper bounds of the harmonic measure on
Theorem 4.4 (Gradient estimates of Green functions).
We have
If then
If then there exists a constant such that
holds for all sufficiently large
@@ 4.2.1. Estimates of harmonic measures
If then is just the geodesic disc centered at with radius It means that for Thus, it follows from (12) that
where is the inward normal derivative on Hence, holds.
Next, we show that holds. Note that is continuous outside and the cut locus of has measure zero, one may assume without loss of generality that Since and as by L’Hôspital’s rule, then it yields from Lemma 4.3 that
for where stands for the inward normal derivative on On the other hand, Lemma 4.3 leads to
Hence, we see that approaches increasingly a geodesic ball centered at with radius as This also implies that
Combining the above, we obtain
Therefore, there exists a constant such that
holds for all sufficiently large ∎
Corollary 4.5 (Estimates of harmonic measures).
We have
If then
If then
for all sufficiently large In the above, is the Riemannian volume element of
@@
The corollary follows immediately from the following relation
4.2.2. Calculus Lemma
In order to establish the calculus lemma which plays an important role in the proof of the second main theorem later, we need to estimate the bounds of the Green function for in terms of integral forms. Set
Lemma 4.6 (Estimates of Green functions).
We have
If then
If then for any there exists sufficiently large such that for we have
Proof.
Theorem 4.7 (Calculus Lemma).
Let be a locally integrable function on Assume that is locally bounded at Then for any there exist a positive constant and a subset with finite Lebesgue measure such that
@@ 4.2.2. Calculus Lemma
If then is the geodesic disc centered at with radius It yields from Theorem 4.6 that
@@
Then
which yields from Corollary 4.5 that
On the other hand, applying Borel lemma to twice, then for any there exists a subset with finite Lebesgue measure such that
holds for all outside Combining (15) with (16), we have
holds for all outside where is a constant.
Next, we show the case where Invoking Lemma 4.6, for any there exists sufficiently large such that for we have
Notice that as due to Lemma 4.3, then there exists sufficiently large such that for we have
Hence, we conclude that
@@
Set
A direct computation leads to
Furthermore, we have
According to Corollary 4.5, there exists sufficiently large such that
holds for all Thus, it follows from (18) that
Using Borel lemma twice, for any , we have
holds for all outside a subset with finite Lebesgue measure. Applying (19) and (20) and noting that as we see that there exists sufficiently large such that for all we have
Therefore, by this with (4.2.2)
holds for all outside where
is a constant. This proves the theorem. ∎
4.2.3. Logarithmic Derivative Lemma
To establish the second main theorem, we still need to prove a logarithmic derivative lemma. Let be the gradient operator on associated with the Kähler metric Let be a meromorphic function on The norm of the gradient of is defined by
in a local holomorphic coordinate where is the inverse of Define
where
Indeed, define for
where
It is not difficult to check that
On take a singular metric
which is satisfied with
To establish the Logarithmic Derivative Lemma, we require the following two lemmas.
@@ 4.2.3. Logarithmic Derivative Lemma
A direct computation leads to
which yields from Fubini’s theorem that
@@
Assume that Then for any there exists a subset with finite Lebesgue measure such that
holds for all outside
@@
Make use of the concavity of , we have
Thus, it yields from Theorem 4.7 that
for all outside The lemma is proved. ∎
@@
Let be a nonconstant meromorphic function on Then for any there exists a subset with finite Lebesgue measure such that
holds for all outside
@@
By Lemma 4.9, for any there exists a subset with finite Lebesgue measure such that
@@ 4.2.4. Second Main Theorem
Let be a nonconstant meromorphic function. Let be distinct values in Then for any there exists a subset with finite Lebesgue measure such that
holds for all outside
@@
in the sense of currents, in which is the zero divisor of and is the ramification divisor of By Dynkin formula (see Lemma 4.1), it yields that
On the other hand, there exist a finite open overing of and rational functions on such that is holomorphic on on and with On each we write
@@
Fix a we can take a local holomorphic coordinate near and a local holomorphic coordinate near such that
@@
By this with Dynkin formula (see Lemma 4.1) and Theorem 4.10, for any there is a subset with finite Lebesgue measure such that
holds for all outside By (4.2.4) and (23), then the theorem holds. ∎
Let us consider the defect relation. Define the defect and the simple defect of with respect to respectively by
According to Theorem 4.2, we obtain Since has maximal volume growth, then whenever is nonconstant, as
Corollary 4.12 (Defect relation).
Assume the same conditions as in Theorem Then
@@ 4.3. Second Main Theorem for non-positive sectional curvature
In the subsection, we assume that has non-positive sectional curvature. Fix a point Review that denotes the geodesic ball centered at with radius By Sard’s lemma, the boundary of is a submanifold of for almost every Set
where is the pointwise lower bound of Ricci curvature defined by
It remains to use to denote the positive Green function of for with a pole at satisfying Dirichlet boundary condition, and use to denote the harmonic measure on with respect to Let be a meromorphic function. Similarly, the Nevanlinna’s functions are defined by
Refer to Dong [14] (see Atsuji [3] also), we have the Second Main Theorem as follows:

### v9

@@ 2.2. Multi-leaf Complex Manifolds of Algebroid Functions
Let be a function element. A neighborhood of is defined as a set of all such that is a direct analytic continuation of satisfying i.e., Taking a proper if necessary, there exists a biholomorphic mapping such that where is the unit ball centered at the coordinate origin in Define a natural mapping
@@
for some two single-valued components of respectively, where is the branching set of for Set
@@
If then it corresponds to meromorphic function elements in In particular, if then reduce to regular function elements. If then it can correspond to multi-leaf function elements in where is a -order branching point of satisfying In other words, can be separated into single-valued components for
@@
Let be single-valued connected components of Define a mapping by
where Note that is a homeomorphism and its inverse is
@@
We will show that is a complex atlas of Suppose It is sufficient to prove the transition mapping
is a biholomorphic mapping. This is automatically a homeomorphism. Via a similar discussion, one only needs to deal with the situation where is a -leaf meromorphic function element and is a regular function element. For every there exists a single-valued component of such that in some neighborhood of i.e., It is clear that because all function elements in are meromorphic. A direct computation leads to
@@
Finally, we treat the singular part Let be the natural projection by mapping a function element into Then, we note that is a generic -sheeted analytic covering such that
Since is an analytic set of complex dimension not greater than then the complex structure of can be viewed as a natural extension of the complex structure of This implies that we can naturally extend the complex structure of to via To conclude, we obtain:
@@
is a non-compact -leaf complex manifold of complex dimension which is called the complex manifold of
@@ 2.3. Zero and Pole Divisors of Algebroid Functions
Let be a -valued algebroid function on defined by equation Let be single-valued meromorphic components of As a matter of convenience, we write
@@
Let be a zero of Since is an empty set or an analytic set of complex codimension 2, then we have in the sense of divisors. Thus, we assume that It yields that
@@
Combining this with (10), we see that reaches the maximal order of a pole of Since then we obtain
which follows that ∎
@@
3. Uniformization of Algebroid Functions
This section works on the uniformization of algebroid functions on a non-compact Hermitian manifold. As shown in Section 2.2, we have constructed a multi-leaf complex manifold of a multi-valued algebroid function on a non-compact Hermitian manifold In fact, this realizes the uniformization of by liftting to a meromorphic function on via the natural projection such that In what follows, we shall provide another way towards the uniformization of algebroid functions without constructing a multi-leaf complex manifold. Our main idea comes from the excellent work of W. Stoll [35] on algebroid reduction.
@@ 3.2. Uniformization
which yields that If then (12) defines an algebraic equation
@@
By this with (11), we have To conclude, we obtain
@@
Set with a basis By (14), we can write
@@
in the sense of isomorphisms. On the other hand, Lemma 3.6 says that there exist a non-compact complex manifold of complex dimension a proper surjective holomorphic mapping of generic sheet number and a holomorphic mapping which separates the fibers of such that is the algebroid reduction of via Hence, we have
@@
Treat the -leaf complex manifold of a -valued algebroid function on we have the natural projection of generic sheet number via which is a proper surjective holomorphic mapping.
@@
4. Nevanlinna Theory of Meromorphic Functions
Let be a non-compact complete Kähler manifold of with Laplace-Beltrami operator The Kähler form of is defined by
in local holomorphic coordinates Moreover, we define the Ricci form of by
@@ 4.1. First Main Theorem
where is the Dirac’s delta function on with a pole at The harmonic measure on with respect to is defined by
@@
Next, we introduce Nevanlinna’s functions. Let be a meormorphic function. Denote by the Fubini-Study form on where is the homogeneous coordinate of with and
@@
Lemma 4.1 (Dynkin formula, [1, 11, 12]).
@@ 4.2. Second Main Theorem for Non-negative Ricci Curvature
Assume that is non-parabolic with non-negative Ricci Curvature. Fix a reference point and denote by the geodesic ball centered at with radius in Denote by the Riemannian volume of The non-parabolicity of means that
Thus, we have the unique minimal positive global Green function of for which can be written as
where is the heat kernel of on Let be the Riemannian distance function of from By Li-Yau’s estimate [18], there exist constants such that
holds for all
4.2.1. Construction of and Nevanlinna’s functions
For define
Since and then we conclude that is a precompact domain containing satisfying that and In fact, the family exhausts i.e., for any sequence with we have
So, the boundary of can be formulated as
By Sard’s theorem, is a submanifold of for almost all
Set
Note that defines the Green function of for with a pole at satisfying Dirichelet boundary condition, i.e.,
Let denote the harmonic measure on with respect to defined by
where is the inward normal derivative on is the Riemannian area element of
Let be a meromorphic function. Take in Section 4.1, we have the Nevanlinna’s functions of :
4.2.2. Gradient estimates of Green functions
@@
holds for all where is given by
@@
According to the definition of Green function for it is immediate that for
Since for then we obtain
Let denote the gradient operator on any Riemannian manifold. Cheng-Yau [8] proved the following theorem.
Lemma 4.4.
Let be a complete Riemannian manifold. Let be a geodesic ball centered at with radius Then, there exists a constant depending only on the dimension of such that
holds for any non-negative harmonic function on where is the Riemannian distance between and and is the lower bound of Ricci curvature of
Theorem 4.5.
There exists a constant such that
@@
By
we obtain an estimate for upper bounds of as follows.
Corollary 4.6.
There exists a constant such that
where is the Riemannian area element of
4.2.3. Calculus Lemma and Logarithmic Derivative Lemma
Lemma 4.7 (Borel Lemma).
Let be a non-decreasing function on with Then for any there exists a subset of finite Lebesgue measure such that such that
@@ 4.2.3. Calculus Lemma and Logarithmic Derivative Lemma
The conclusion is clearly true for Next, we assume that Since is a non-decreasing function, then there exists a number such that The non-decreasing property of implies that the limit exists or equals If then Set
Note that exists for almost all Then, we have
This completes the proof. ∎
@@
Theorem 4.8 (Calculus Lemma).
Let be a locally integrable function on Assume that is locally bounded at Then there exists a constant such that for any there exists a subset of finite Lebesgue measure such that
holds for all outside where is given by
Proof.
Invoking Lemma 4.3, we have
Set
A simple computation leads to
In further, we have
Apply Borel’s lemma to the left hand side of the above equality twice: one is to and the other is to then we deduce that for any there exists a subset of finite Lebesgue measure such that
holds for all outside On the other hand, Corollary 4.6 implies that there exists a constant such that
Set Combining the above, we have
holds for all outside ∎
In the following, we establish a logarithmic derivative lemma. Let be a meromorphic function on The norm of the gradient of is defined by
in local holomorphic coordinates here is the inverse of Define the Nevanlinna’s characteristic function of by
@@
It is not difficult to show that
On one puts a singular metric
which leads to
@@
Note that
Whence, it concludes from Fubini’s theorem that
@@
Let be a meromorphic function on Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside where is given by
@@
Let be a meromorphic function on Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside where is given by
@@
By Lemma 4.18, we have for any there exists a subset of finite Lebesgue measure such that
@@ 4.2.4. Second Main Theorem
Let be a nonconstant meromorphic function. Let be distinct values in Then for any there exists a subset of finite Lebesgue measure such that
holds for all outside where is given by
@@
in the sense of currents, in which is the zero divisor of and is the ramification divisor of By Dynkin formula (see Lemma 4.1)
On the other hand, there exist a finite open overing of and rational functions on such that is holomorphic on on and
for On each write
@@
For any we can take local holomorphic coordinates near and a local holomorphic coordinate near such that
@@
By this with Dynkin formula and Theorem 4.11, for any there exists a subset of finite Lebesgue measure such that
holds for all outside where one used the fact that Combining (4.2.4) and (19), we have the theorem proved. ∎
@@ 4.3. Second Main Theorem for Non-positive Sectional Curvature
Assume that has non-positive sectional curvature satisfying (3). From the definition of curvatures, it is not difficult to see that In particular, we have if has constant sectional curvature Fix a reference point We may assume without loss of generality that is a pole of for a technical reason if necessary, since, otherwise, one can consider the universal covering As long as is equipped with the pull-back metric induced from via then has the same curvatures as ones of In fact, is a Cartan-Hadamard manifold. In the following, we denote by the Riemannian distance function of from on
4.3.1. Construction of and Nevanlinna’s functions
A. Setups for
Due to the curvature conditions, we have a unique minimal positive global Green function of with a pole at for Consider the following initial value problems on :
and
Evidently, the two ODEs have the unique solutions
Using the estimates of we have (see [17, 34])
where
Here, is the standard Euclidean area of For define
Since as and then we deduce that is a precompact domain containing in which exhausts So, the boundary of can be formulated as
Set
which defines the Green function of for with a pole at satisfying Dirichlet boundary condition. Denote by the harmonic measure on with respect to i.e.,
where is the inward normal derivative on is the Riemannian area element of In particular, if In this case, we have
which is just the geodesic ball centered at with radius in Then
Let be the geodesic ball centered at with radius in We give the relation between and as follows.
Set
Theorem 4.13.
For we have and
Hence, there exists a constant independent of such that
Proof.
B. Setups for
We shall extend the settings for to the case that For the function does not make sense. However, since
we can extend continuously to defined by (4).
Notice that is parabolic when i.e., there exist no positive global Green functions for In other words, for Hence, one cannot use to define the domain in this case. However, we note (21) for it is thus natural to define
for Let it then yields from (22) that
To conclude, has the unified form
for In further, we have
for
Let be a meromorphic function. Similarly, one can defined the Nevanlinna’s functions and of as well as on as in Section 4.2.1 and Section 4.2.4.
4.3.2. Calculus Lemma and Logarithmic Derivative Lemma
Lemma 4.14.
For we have for
Proof.
When is the geodesic ball centered at with radius The conclusion holds immediately due to (23). In the following, we consider the other case. Since on then for
This completes the proof. ∎
Theorem 4.15.
There exists a constant such that
holds for
Proof.
When is a geodesic ball centered at with radius Hence, it yields from (4) and (23) that
This implies the conclusion holds. For by Lemma 4.4 with (3), (4) and (20) (see Remark 5 in [34] also)
for some large constant which depends only on the dimension By (20) again
which gives for So, we conclude that
where ∎
Corollary 4.16.
There exists a constant such that
holds for where is the Riemannian area element of
Set
Theorem 4.17 (Calculus Lemma).
Let be a locally integrable function on Assume that is locally bounded at Then for any there exists a set of finite Lebesgue measure such that
holds for outside where is defined by
Proof.
Lemma 4.14 yields that
Differentiating to obtain
In further
Applying Borel’s lemma to the left hand side on the above equality twice, then we see that there exists a set of finite Lebesgue measure such that
holds for outside Meanwhile, Corollary 4.16 means that there exists a constant such that
Substituting (27) into (4.3.2), we conclude that
for outside , where ∎
Lemma 4.18.
Let be a meromorphic function on Then for any there exists a subset of finite Lebesgue measure such that
holds for outside where is defined by
Proof.
Theorem 4.19 (Logarithmic Derivative Lemma).
Let be a meromorphic function on Then for any there exists a subset of finite Lebesgue measure such that
holds for outside where is defined by
Proof.
4.3.3. Second Main Theorem